OpenStudy (anonymous):

not really a maths question, but how do you write fractions in the equation window in the post box? (and not just using /)

OpenStudy (angela210793):

i don't know that too....

OpenStudy (anonymous):

there's no need to use the equation editor if you learn how to use LaTeX

OpenStudy (anonymous):

I'm not entirely sure. A lot of people just copy and paste from a website called WolfRamAlpha.com

OpenStudy (angela210793):

u can't cope-paste from Wolfram

OpenStudy (anonymous):

\frac{numerator}{denominator} that's what you can use to write fractions. To render it, enclose it between square brackets escaped by backslashes so it looks like \[\frac{numerator}{denominator}\]

OpenStudy (anonymous):

or if you want it to be on the same line, use parentheses instead of square brackets: \( \frac{3}{2}\)

OpenStudy (anonymous):

oh right, thanks. I just find it neater. thats really helpful

OpenStudy (anonymous):

\begin{array}l\color{#FF0000}{\text{LaTex}}\text{ }\color{#FF7F00}{\text{is}}\text{ }\color{#FFE600}{\text{great}}\text{ }\color{#00FF00}{\text{to}}\text{ }\color{#0000FF}{\text{learn!}}\end{array}

OpenStudy (jagatuba):

Thanks agdgdgdgwngo. I was not aware of the parenthesis trick.